OpenStudy (anonymous):

Using the substition method find the ordered pairs for 5a = 3-b and 5a = 4 -b

OpenStudy (mertsj):

Use the transitive property. If a = b and b = c than a = c

OpenStudy (mertsj):

If 3-b=5a and 5a = 4-b then 3-b=4-b

OpenStudy (shaik0124):

but this will have no solution

OpenStudy (shaik0124):

i think there is some thing error in ur question either it must 4+b or 3+b

OpenStudy (anonymous):

I thought it had no solution but wasn't sure wanted to bounced it off someone's else. No the second part of the question was did it have an ordered pair, no solution or infinate. I chose no solution

OpenStudy (shaik0124):

no solutionis answer

OpenStudy (anonymous):

thanks that is what I thought